OpenStudy (anonymous):

Solve each equation on the interval [0,2π): a. 4sin2x – 3 = 0 b. cos(3x) = -1

OpenStudy (anonymous):

@freckles @IrishBoy123 @ganeshie8

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@Jennjuniper

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

idk how to do this hun. sorry.

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

@e.mccormick is offline btw

OpenStudy (anonymous):

@peachpi I'm assuming you don't know how to do this either??

OpenStudy (anonymous):

@TheSmartOne @whpalmer4

OpenStudy (irishboy123):

a. \(4sin2x – 3 = 0\) thus: \(4sin2x = 3\) you can finish that off with your calculator. solve for '2x' first.

OpenStudy (anonymous):

Wait, so do I solve 4sin(2x) with my calculator? What about the " = 3" that you put. Should the answer to 4sin(2x) be my answer to "a"? What about b? @IrishBoy123

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@GenTorr

OpenStudy (anonymous):

can yo u help with my questinon

OpenStudy (gentorr):

i can try

OpenStudy (anonymous):

@logan6767 depends on what your question is...

OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4

OpenStudy (anonymous):

(1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

Sorry, can't solve it @logan6767

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

4 sin (2x) - 3 = 0 4 sin (2x) = 3 sin (2x) = 3/4 2x = arcsin (3/4) → this is the 1st quadrant solution for 2x x = ½ arcsin (3/4) 2x = π - arcsin (3/4) → this is the 2nd quadrant solution for 2x x = ½(π - arcsin (3/4)) If 0<x<2π, then 0<2x<4π, so other solutions for 2x are 2x = 2π + arcsin (3/4) x = π + ½ arcsin (3/4) 2x = 3π - arcsin (3/4) x = ½(3π - arcsin (3/4))

OpenStudy (anonymous):

Is this correct or are you guessing??

OpenStudy (anonymous):

and I'm out

OpenStudy (anonymous):

what about b?

OpenStudy (anonymous):

So a has 6 solutions?: 2x = 2π + arcsin (3/4) x = π + ½ arcsin (3/4) 2x = 3π - arcsin (3/4) x = ½(3π - arcsin (3/4)) x = ½ arcsin (3/4) x = ½(π - arcsin (3/4)) Or 4 solutions?: x = π + ½ arcsin (3/4) x = ½(3π - arcsin (3/4)) x = ½ arcsin (3/4) x = ½(π - arcsin (3/4)) @peachpi

OpenStudy (anonymous):

4

OpenStudy (anonymous):

okay, but what about b??

OpenStudy (anonymous):

The angle here is 3x, so the interval is [0, 6π). Pull out your unit circle and find the angle that has a cosine of -1. Then add 2π twice to get the other 2 solutions

OpenStudy (anonymous):

Do I add 2pi to the -1 or to the angle that has a cosine of -1?